# Double-coset zeta functions for groups acting on trees

### v1

@@ 3.4.2. The universal group $U(\Delta,\mathbb{T})$
In other words, $U(\Delta,\mathbb{T})$ collects all elements of $\mathrm{Aut}_{\pi}(T)$ acting on $o^{-1}(v)$ as a permutation in $G(\pi(v))$, for every $v\in VT$. If $\mathbb{T}$ is the standard $\Delta$-tree associated to $\iota$ and $c_{0}$ (cf. Section 3.4.1), we write $U(\Delta,\iota,c_{0})$ instead of $U(\Delta,\mathbb{T})$. If $T$ is locally finite, the group $U(\Delta,\mathbb{T})$ is a t.d.l.c. group with respect to the subspace topology from $\mathrm{Aut}(T)$. Indeed, $U(\Delta,\mathbbm{T})$ is a closed subgroup of the t.d.l.c. group $\mathrm{Aut}(T)$ (cf. [17, §6]). If additionally $\Gamma$ is finite, then $U(\Delta,\mathbbm{T})$ is a compactly generated t.d.l.c. group (cf. [17, Proposition 6.5]).
@@
      Set $F:=G(c)$ and assume that $G(d)=C_{2}$, $|X_{c}|=k\geq 2$ and $|X_{d}|=2$. Adapting [18, Example 11] to actions on trees without edge inversion, the action $(U(\Delta,\mathbb{T}),T)$ is permutational isomorphic to the action of the Burger–Mozes universal group $U(F)$ on the barycentric subdivision $T_{k}^{\prime}$ of the $k$-regular tree $T_{k}$ (cf. [5, §3.2]). Here we consider $T_{k}^{\prime}$, and not $T_{k}$, because $U(F)$ acts vertex-transitively (and thus with edge inversions) on $T_{k}$.
@@
      Set $F_{1}:=G(c)$ and $F_{2}:=G(d)$. Following [18, Example 12], the action $U(\Delta,\mathbb{T})$-action on $T$ is permutational isomorphic to $(U(F_{1},F_{2}),T)$, where $U(F_{1},F_{2})$ is the group introduced by S. Smith in [22].
The following fact collects some key results of the work of C. Reid and S. Smith [17]. It motivates why, throughout the paper, we focus on actions of the form $(U(\Delta,\iota,c_{0}),T(\Delta,\iota,c_{0}))$ while considering (P)-closed actions on trees with associated local action diagram $\Delta$.
@@
    ([17, Lemma 3.5, Theorem 3.12]) Let $\Delta$ be a local action diagram. For all $\Delta$-trees $\mathbb{T}=(T,\pi,\mathcal{L})$ and $\mathbb{T}^{\prime}=(T^{\prime},\pi^{\prime},\mathcal{L}^{\prime})$ and $v\in VT$, $v^{\prime}\in VT^{\prime}$ with $\pi(v)=\pi^{\prime}(v^{\prime})$, there is a graph isomorphism $\phi\colon T\to T^{\prime}$ such that $\phi(v)=v^{\prime}$, $\pi=\pi^{\prime}\circ\phi$ and $\phi U(\Delta,\mathbbm{T})\phi^{-1}=U(\Delta,\mathbb{T}^{\prime})$.
@@
    ([17, Theorem 3.9]) Let $\Delta$ be a local action diagram and $\mathbb{T}=(T,\pi,\mathcal{L})$ be a $\Delta$-tree. Then the local action diagram associated to $(U(\Delta,\mathbb{T}),T)$ is isomorphic to $\Delta$ (in the sense of [17, Definition 3.2]).
@@
    ([17, Theorem 3.10]) Let $(G,T)$ be a (P)-closed action on a tree with associated local action diagram $\Delta$. Then $(G,T)$ is permutational isomorphic to $(U(\Delta,\iota,c_{0}),T(\Delta,\iota,c_{0}))$, for every inversion $\iota$ in $\Delta$ and every $c_{0}\in V\Gamma$.
@@ 3.5. Weakly locally $\infty$-transitive actions on trees
Let $(G,T)$ be a group action on a tree with quotient map $\pi\colon T\to\Gamma=G\backslash T$. Then $(G,T)$ is said to be weakly locally $\infty$-transitive at $v\in VT$ if, for every path $\mathfrak{p}$ in $\Gamma$ starting at $\pi(v)$, the stabiliser $G_{v}$ acts transitively on the set
@@
Let $\pi\colon T\to\Gamma=G\backslash T$ be the quotient map of $(G,T)$. Given $v\in VT$ with $\pi(v)=c$ and for every $d\geq 0$, we observe that
(3.12) $\{\tilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(v\to T)\mid\ell(\tilde{\mathfrak{p}})=d\}=\bigsqcup_{\mathfrak{p}\in\mathcal{P}_{\Gamma}(c\to\Gamma),\atop\ell(\mathfrak{p})=d}\{\tilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(v\to T)\mid\pi(\tilde{\mathfrak{p}})=\mathfrak{p}\}.$
Note that $\Gamma$ is a $1$-segment if, and only if, |$|\{\mathfrak{p}\in\mathcal{P}_{\Gamma}(c\to\Gamma)\,:\,\ell(\mathfrak{p})=d\}|=1$ for all $d\geq 0$ and $c\in V\Gamma$. If $(G,T)$ is weakly locally $\infty$-transitive and $\Gamma$ is a $1$-segment, then $G_{v}$ acts transitively on $\{\tilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(v\to T)\mid\ell(\tilde{\mathfrak{p}})=d\}$, for all $v\in VT$ and $d\geq 0$. Conversely, if $(G,T)$ is locally $\infty$-transitive, then $(G,T)$ is edge-transitive. More precisely, since $G_{v}$ acts transitively on $o^{-1}(v)$ for every $v\in VT$, we deduce that $\Gamma$ is a $1$-segment. Indeed, if $\Gamma$ is a $1$-loop, for every $v\in VT$ the group $G_{v}$ has two orbits on $o^{-1}(v)$. Moreover, for all $v\in VT$ and $d\geq 0$, there is a unique path $\mathfrak{p}$ in $\Gamma$ starting at $\pi(v)$ of length $d$. By (3.12), we conclude that $(G,T)$ is weakly locally $\infty$-transitive. ∎
@@
    for all $a,b\in o^{-1}(c)$ and $x\in X_{a}$, the group $G(c)_{x}$ acts transitively on $X_{b}\setminus\{x\}$.
Conversely, if $(G,T)$ is (P)-closed, condition ($\diamondsuit$) implies that $(G,T)$ is weakly locally $\infty$-transitive at $v$.
@@
To prove the first part of the assertion, let $a,b\in o^{-1}(c)$. By essential uniqueness of the local action diagram associated to $(G,T)$ (cf. Section 3.4), we may assume that $X_{a},X_{b}\subseteq o^{-1}(v)$. Since $(G,T)$ is weakly locally $\infty$-transitive at $v$, for every $e\in X_{a}$ the stabiliser $G_{e}$ acts transitively on
@@
Let now $(G,T)$ be (P)-closed and suppose that ($\diamondsuit$) holds. Without loss of generality, set $G=U(\Delta,\iota,c)$ and $T=T(\Delta,\iota,c)$. Consider two geodesics in $T$, say $[v,w_{1}]=(e_{1},\dots,e_{n})$ and $[v,w_{2}]=(f_{1},\dots,f_{n})$ for some $n\geq 1$ and $w_{1},w_{2}\in VT$, with the same image in $\Gamma$. For every $1\leq i\leq n$, since $\pi(e_{i})=\pi(f_{i})$, there exists $g_{i}\in G$ such that $f_{i}=g_{i}\cdot e_{i}$. Note that $o(g_{i}\cdot e_{i+1})=g_{i}\cdot t(e_{i})=t(f_{i})=o(f_{i+1})$ and $\pi(g_{i}\cdot e_{i+1})=\pi(e_{i+1})=\pi(f_{i+1})$. Hence, both $\mathcal{L}(g_{i}\cdot e_{i+1})$ and $\mathcal{L}(f_{i+1})$ belong to $X_{\pi(f_{i+1})}$. By ($\diamondsuit$) applied to the vertex $c_{i}=o(\pi(f_{i+1}))$, for every $i<n$ there is $h_{i}\in G_{\bar{f}_{i}}$ such that $f_{i+1}=h_{i}g_{i}\cdot e_{i+1}$. For every $i<n$, set $k_{i}:=h_{i}g_{i}$ and observe that
@@
In particular, $k_{1}\in G_{v}$ and $k_{i}\cdot e_{i+1}=f_{i+1}=k_{i+1}\cdot e_{i+1}$ for every $i<n-1$. For every $i<n-1$ write $u_{i+1}:=k_{i}^{-1}k_{i+1}\in G_{e_{i+1}}$, and let $u_{i+1}^{-}\in G_{T_{\geq\bar{e}_{i+1}}}$ and $u_{i+1}^{+}\in G_{T_{\geq e_{i+1}}}$ be such that $u_{i+1}=u_{i+1}^{-}u_{i+1}^{+}$ (recall that $(G,T)$ is (P)-closed). Set $\tilde{k}_{1}:=k_{1}$ and, for every $12\leq i\leq n-$ 1, define inductively $\tilde{k}_{i}:=\tilde{k}_{i-1}u_{i}^{-}$. For each $12\leq i\leq n-$ 1, note that $k_{i}=\tilde{k}_{i}u_{i}^{+}$ and, since $u_{i}^{+}$ fixes both $e_{i}$ and $e_{i+1}$, from (3.13) we have
@@
Moreover, we claim that $\tilde{k}_{i}$ fixes $v$ for every $11\leq i\leq n-$ 1. If $i=1$, this is clear. For $i\geq 2$, assuming inductively that $\tilde{k}_{i-1}$ fixes $v$, the fact that $u_{i}^{-}$ fixes $T_{\geq\bar{e}_{i}}$ pointwise implies that $\tilde{k}_{i}=\tilde{k}_{i-1}u_{i}^{-}$ fixes $v$. In particular, $\tilde{k}_{n-1}\in G_{v}$ and (3.14) yields
@@
Being $T$ a tree, we conclude that $[v,w_{2}]=[v,\tilde{k}_{n-1}\cdot w_{1}]=\tilde{k}_{n}\cdot[v,w_{1}]$. ∎
@@
    Let $(T,\pi)$ be the universal cover of a connected edge-weighted graph $(\Gamma,\omega)$ as in Example 1.1. According to Example 3.7(i), one checks that ($\diamondsuit$) is satisfied for every $c\in V\Gamma$. Hence, by Proposition 3.11, $(\mathrm{Aut}_{\pi}(T),T)$ is weakly locally $\infty$-transitive. This gives an alternative proof of [7, Theorem 3.1 and the comment thereafter].
@@
    Let $U(F_{1},F_{2})$ be the Smith’s group associated to two transitive groups $F_{1}\leq\mathrm{Sym}(\{1,\ldots,k_{1}\})$ and $F_{2}\leq\mathrm{Sym}(\{1,\ldots,k_{2}\})$, $k_{1},k_{2}\geq 2$. According to Example 3.7(iib), condition ()($\diamondsuit$)( ♢ ) is satisfied for every $c\in V\Gamma$ if, and only if, both $F_{1}$ and $F_{2}$ are $2$-transitive. Moreover, the $U(F_{1},F_{2})$-action on the $(k_{1},k_{2})$-biregular tree $T_{k_{1},k_{2}}$ has quotient graph a $1$-segment. Hence, as in (ii), $(U(F_{1},F_{2}),T_{k_{1},k_{2}})$ is weakly locally $\infty$-transitive if, and only if, it is locally $\infty$-transitive. By Proposition 3.11 we conclude that $(U(F_{1},F_{2}),T_{k_{1},k_{2}})$ is locally $\infty$-transitive if, and only if, both $F_{1}$ and $F_{2}$ are $2$-transitive.

### v2

@@ 3.4.2. The universal group $U(\Delta,\mathbb{T})$
In other words, $U(\Delta,\mathbb{T})$ collects all elements of $\mathrm{Aut}_{\pi}(T)$ acting on $o^{-1}(v)$ as a permutation in $G(\pi(v))$, for every $v\in VT$.
If $\mathbb{T}$ is the standard $\Delta$-tree associated to $\iota$ and $c_{0}$ (cf. Section 3.4.1), we write $U(\Delta,\iota,c_{0})$ instead of $U(\Delta,\mathbb{T})$.
If $T$ is locally finite, the group $U(\Delta,\mathbb{T})$ is a t.d.l.c. group with respect to the subspace topology from $\mathrm{Aut}(T)$. Indeed, $U(\Delta,\mathbbm{T})$ is a closed subgroup of the t.d.l.c. group $\mathrm{Aut}(T)$ (cf. [24, §6]). If additionally $\Gamma$ is finite, then $U(\Delta,\mathbbm{T})$ is a compactly generated t.d.l.c. group (cf. [24, Proposition 6.5]).
@@
      Set $F:=G(c)$ and assume that $G(d)=C_{2}$, $|X_{c}|=k\geq 2$ and $|X_{d}|=2$. Adapting [23, Example 11] to actions on trees without edge inversion, the action $(U(\Delta,\mathbb{T}),T)$ is permutational isomorphic to the action of the Burger–Mozes universal group $U(F)$ on the barycentric subdivision $T_{k}^{\prime}$ of the $k$-regular tree $T_{k}$ (cf. [5, §3.2]). Here we consider $T_{k}^{\prime}$, and not $T_{k}$, because $U(F)$ acts vertex-transitively (and thus with edge inversions) on $T_{k}$.
@@
      Set $F_{1}:=G(c)$ and $F_{2}:=G(d)$. Following [23, Example 12], the action $U(\Delta,\mathbb{T})$-action on $T$ is permutational isomorphic to $(U(F_{1},F_{2}),T)$, where $U(F_{1},F_{2})$ is the group introduced by S. Smith in [29].
The following theorem collects some key results of the work of C. Reid and S. Smith [24]. It motivates why, throughout the paper, we focus on actions of the form $(U(\Delta,\iota,c_{0}),T(\Delta,\iota,c_{0}))$ while considering (P)-closed actions on trees with associated local action diagram $\Delta$. Note that the result below holds for general group actions on trees and general local action diagrams (even if they do not satisfy Conventions 3.0.1 and 3.1.1, respectively).
@@
    ([24, Lemma 3.5, Theorem 3.12]) Let $\Delta$ be a local action diagram. For all $\Delta$-trees $\mathbb{T}=(T,\pi,\mathcal{L})$ and $\mathbb{T}^{\prime}=(T^{\prime},\pi^{\prime},\mathcal{L}^{\prime})$ and $v\in VT$, $v^{\prime}\in VT^{\prime}$ with $\pi(v)=\pi^{\prime}(v^{\prime})$, there is a graph isomorphism $\phi\colon T\to T^{\prime}$ such that $\phi(v)=v^{\prime}$, $\pi=\pi^{\prime}\circ\phi$ and $\phi U(\Delta,\mathbbm{T})\phi^{-1}=U(\Delta,\mathbb{T}^{\prime})$.
@@
    ([24, Theorem 3.9]) Let $\Delta$ be a local action diagram and $\mathbb{T}=(T,\pi,\mathcal{L})$ be a $\Delta$-tree. Then the local action diagram associated to $(U(\Delta,\mathbb{T}),T)$ is isomorphic to $\Delta$ (in the sense of [24, Definition 3.2]).
@@
    ([24, Theorem 3.10]) Let $(G,T)$ be a (P)-closed action on a tree with associated local action diagram $\Delta$. Then $(G,T)$ is permutational isomorphic to $(U(\Delta,\iota,c_{0}),T(\Delta,\iota,c_{0}))$, for every inversion $\iota$ in $\Delta$ and every $c_{0}\in V\Gamma$.
@@ 3.5. Weakly locally $\infty$-transitive actions on trees
Recall that group actions on trees satisfy Convention 3.0.1 and local action diagrams satisfy Convention 3.1.1.
@@
Let $(G,T)$ be a group action on a tree with quotient map $\T\pi\colon T\to\Gamma=G\backslash$ T. Then $(G,T)$ is said to be weakly locally $\infty$-transitive at $v\in VT$ if, for every path $\mathfrak{p}$ in $\Gamma$ starting at $\pi(v)$, the stabiliser $G_{v}$ acts transitively on the set
@@
Let $\T\pi\colon T\to\Gamma=G\backslash$ T be the quotient map of $(G,T)$. Given $v\in VT$ with $\pi(v)=c$ and for every $d\geq 0$, we observe that
(3.12) $\{\tilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(v\to T)\mid\ell(\tilde{\mathfrak{p}})=d\}=\hskip-5.69046pt\bigsqcup_{\mathfrak{p}\in\mathcal{P}_{\Gamma}(c\to\Gamma),\atop\ell(\mathfrak{p})=d}\hskip-5.69046pt\{\tilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(v\to T)\mid\pi(\tilde{\mathfrak{p}})=\mathfrak{p}\}.$
Note that $\Gamma$ is a $1$-segment if, and only if, $|\{\mathfrak{p}\in\mathcal{P}_{\Gamma}(c\to\Gamma)\,:\,\ell(\mathfrak{p})=d\}|=1$ for all $d\geq 0$ and $c\in V\Gamma$.
If $(G,T)$ is weakly locally $\infty$-transitive and $\Gamma$ is a $1$-segment, then $G_{v}$ acts transitively on $\{\tilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(v\to T)\mid\ell(\tilde{\mathfrak{p}})=d\}$, for all $v\in VT$ and $d\geq 0$, and $(G,T)$ is locally $\infty$-transitive.
Conversely, if $(G,T)$ is locally $\infty$-transitive, then $(G,T)$ is arc-transitive. More precisely, since $G_{v}$ acts transitively on $o^{-1}(v)$ for every $v\in VT$, we deduce that $\Gamma$ is a $1$-segment. Indeed, if $\Gamma$ is a $1$-loop, for every $v\in VT$ the group $G_{v}$ has two orbits on $o^{-1}(v)$. Finally, for all $v\in VT$ and $d\geq 0$, there is a unique path $\mathfrak{p}$ in $\Gamma$ starting at $\pi(v)$ of length $d$. By (3.12), $(G,T)$ is also weakly locally $\infty$-transitive. ∎
@@
    for all $a,b\in o^{-1}(c)$ and all $x\in X_{a}$, the group $G(c)_{x}$ acts transitively on $X_{b}\setminus\{x\}$.
Conversely, if $(G,T)$ is (P)-closed, having condition $(\diamondsuit)$ for all $c\in V\Gamma$ implies that $(G,T)$ is weakly locally $\infty$-transitive.
@@
To prove the first part of the assertion, let $a,b\in o^{-1}(c)$. By the essential uniqueness of the local action diagram associated to $(G,T)$ (cf. Section 3.4), we may assume that $X_{a},X_{b}\subseteq o^{-1}(v)$. Since $(G,T)$ is weakly locally $\infty$-transitive at $v$, for every $e\in X_{a}$ the stabiliser $G_{e}$ acts transitively on
@@
Let now $(G,T)$ be (P)-closed and suppose that ($\diamondsuit$) holds for all $c\in V\Gamma$. Without loss of generality, let $G=U(\Delta,\iota,c)$, $T=T(\Delta,\iota,c)$ and $v\in VT$ be the root of $T(\Delta,\iota,c)$ (so $\pi(v)=c$). Consider two geodesics in $T$ with the same image in $\Gamma$, say $v,w_{1}]=(e_{1},\dots,e_{n})$ and $v,w_{2}]=(f_{1},\dots,f_{n})$ for some $n\geq 1$ and $w_{1},w_{2}\in VT$. For every $1\leq i\leq n$, since $\pi(e_{i})=\pi(f_{i})$, there exists $g_{i}\in G$ such that $f_{i}=g_{i}\cdot e_{i}$. Note that $o(g_{i}\cdot e_{i+1})=g_{i}\cdot t(e_{i})=t(f_{i})=o(f_{i+1})$ and $\pi(g_{i}\cdot e_{i+1})=\pi(e_{i+1})=\pi(f_{i+1})$. Hence, both $\mathcal{L}(g_{i}\cdot e_{i+1})$ and $\mathcal{L}(f_{i+1})$ belong to $X_{\pi(f_{i+1})}$.
By ($\diamondsuit$) applied to the vertex $c_{i}=o(\pi(f_{i+1}))$, for every $i<n$ there is $h_{i}\in G_{\bar{f}_{i}}$ such that $f_{i+1}=h_{i}g_{i}\cdot e_{i+1}$. For every $i<n$, set $k_{i}:=h_{i}g_{i}$ and observe that
@@
In particular, $k_{1}=h_{1}g_{1}\in G_{v}$ (as $h_{1},g_{1}\in G_{v}$) and $k_{i}\cdot e_{i+1}=f_{i+1}=k_{i+1}\cdot e_{i+1}$ for every $1\leq i<n-1$. From $(k_{i})_{1\leq i\leq n-1}$, we now construct a sequence $(\tilde{k}_{i})_{1\leq i\leq n-1}$ in $G_{v}$ such that $v,w_{2}]=\tilde{k}_{n-1}\cdot[v,w_{1}]$.
For every $1\leq i<n-1$ write $u_{i+1}:=k_{i}^{-1}k_{i+1}\in G_{e_{i+1}}$, and let $u_{i+1}^{-}\in G_{T_{\geq\bar{e}_{i+1}}}$ and $u_{i+1}^{+}\in G_{T_{\geq e_{i+1}}}$ be such that $u_{i+1}=u_{i+1}^{-}u_{i+1}^{+}$ (recall that $(G,T)$ is (P)-closed, so (3.2) holds). For all $2\leq i<j\leq n-1$, $u_{i}^{+}$ has support in $T_{\geq\bar{e}_{i}}$ and $u_{j}^{-}$ has support in $T_{\geq e_{j}}$, which is disjoint to $T_{\geq\bar{e}_{i}}$. Hence, for all $2\leq i<j\leq n-1$ one has
(3.14) $u_{i}^{+}u_{j}^{-}=u_{j}^{-}u_{i}^{+}.$
Let $\tilde{k}_{1}:=k_{1}$ and, for every $2\leq i\leq n-1$, define inductively $\tilde{k}_{i}:=\tilde{k}_{i-1}u_{i}^{-}$. For each $2\leq i\leq n-1$, we claim that
(3.15) $k_{i}=\tilde{k}_{i}u_{2}^{+}u_{3}^{+}\cdots u_{i}^{+}.$
For $i=2$, one has $k_{2}=k_{1}u_{2}=(\tilde{k}_{1}u_{2}^{-})u_{2}^{+}=\tilde{k}_{2}u_{2}^{+}$. For $i\geq 3$, assuming (3.15) true for i1i-1, we have
[…]
Since $u_{2}^{+},\ldots,u_{i}^{+}$ fix both $e_{i}$ and $e_{i+1}$, from (3.13) and (3.15) we have
@@
Moreover, we claim that $\tilde{k}_{i}$ fixes $v$ for every $1\leq i\leq n-1$. If $i=1$, this is clear. For $i\geq 2$, assume inductively that $\tilde{k}_{i-1}$ fixes $v$. Since $u_{i}^{-}$ fixes $T_{\geq\bar{e}_{i}}\ni v$ pointwise, we have that $\tilde{k}_{i}=\tilde{k}_{i-1}u_{i}^{-}$ fixes $v$. In particular, $\tilde{k}_{n-1}\in G_{v}$ and (3.16) yields
@@
Being $T$ a tree, we conclude that $v,w_{2}]=[v,\tilde{k}_{n-1}\cdot w_{1}]=\tilde{k}_{n-1}\cdot[v,w_{1}]$. ∎
@@
    Let $(T,\pi)$ be the universal cover of a connected edge-weighted graph $(\Gamma,\omega)$ as in Example 1.1. According to Example 3.7(i), one checks that ($\diamondsuit$) is satisfied for every $c\in V\Gamma$. Hence, by Proposition 3.11, $(\mathrm{Aut}_{\pi}(T),T)$ is weakly locally $\infty$-transitive. This gives an alternative proof of [9, Theorem 3.1 and the comment thereafter].
@@
    Let $U(F_{1},F_{2})$ be the Smith group associated to two transitive groups $F_{1}\leq\mathrm{Sym}(\{1,\ldots,k_{1}\})$ and $F_{2}\leq\mathrm{Sym}(\{1,\ldots,k_{2}\})$, $k_{1},k_{2}\geq 2$. According to Example 3.7(iib), condition $(\diamondsuit)$ is satisfied for every $c\in V\Gamma$ if, and only if, both $F_{1}$ and $F_{2}$ are $2$-transitive. Moreover, the $U(F_{1},F_{2})$-action on the $(k_{1},k_{2})$-biregular tree $T_{k_{1},k_{2}}$ has quotient graph a $1$-segment. Hence, as in (ii), $(U(F_{1},F_{2}),T_{k_{1},k_{2}})$ is weakly locally $\infty$-transitive if, and only if, it is locally $\infty$-transitive. By Proposition 3.11 we conclude that $(U(F_{1},F_{2}),T_{k_{1},k_{2}})$ is locally $\infty$-transitive if, and only if, both $F_{1}$ and $F_{2}$ are $2$-transitive.
As one may guess, there are (P)-closed actions that are not weakly locally $\infty$-transitive (and possibly not arc-transitive).
Example 3.13.
For $n\geq 3$, recall that the dihedral group $D_{2n}$ of size $2n$ is the group of symmetries of an Euclidean regular $n$-gon, say with vertices $x_{1},\ldots,x_{n}$. Hence, we may consider $D_{2n}$ a subgroup of $\mathrm{Sym}(\{x_{1},\ldots,x_{n}\})$. Note that $D_{2n}$ acts transitively on $\{x_{1},\ldots,x_{n}\}$ (think of the rotations in the correspondent $n$-gon). However, for $n\geq 4$, the $D_{2n}$-action on $\{x_{1},\ldots,x_{n}\}$ is no longer $2$-transitive (as $D_{2n}$ preserves adjacent vertices in the $n$-gon).
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be any local action diagram with the following property: there is $c\in V\Gamma$ such that $o^{-1}(c)=\{a\}$ (so, $c$ is a leaf of $\Gamma$) and, keeping the notation above, such that $G(c)=D_{2n}\leq\mathrm{Sym}(X_{c})$, where $X_{c}=X_{a}=\{x_{1},\ldots,x_{n}\}$ for some $n\geq 4$.
For example, $\Gamma$ can be the graph as depicted below.
[…]
Moreover, one can take $X_{a}=X_{\bar{b}}:=\{x_{1},\ldots,x_{n}\}$, $n\geq 4$, $G(c)=G(e):=D_{2n}$ as before, and let $G(d)$ be any permutation group acting on a finite set $X_{d}$ with two distinct orbits, say $X_{\bar{a}}$ and $X_{b}$ (e.g., $G(d)=2\mathbb{Z}/4\mathbb{Z}$ acting on $X_{d}=\mathbb{Z}/4\mathbb{Z}$ by left translation).
If $\mathbb{T}=(T,\pi,\mathcal{L})$ is any $\Delta$-tree, the action $(U(\Delta,\mathbb{T}),T)$ is clearly (P)-closed but, by Proposition 3.11, not weakly locally $\infty$-transitive. Moreover, unless $\Gamma$ is a $1$-segment, the latter action is also not arc-transitive.
There are also (weakly) locally $\infty$-transitive group actions $(G,T)$ on trees that are not (P)-closed but $G\leq\mathrm{Aut}(T)$ is closed. The following example is based on [14, Example 6.33].
Example 3.14.
Let $T=T_{p+1}$ be the $(p+1)$-regular tree with $p$ prime. Recall by Example 3.3(ii) that $\mathrm{PSL}_{2}(\mathbb{Q}_{p})$ acts faithfully on $T$, so it can be embedded in $\mathrm{Aut}(T)$. Let $G$ be the closure of $\mathrm{PSL}_{2}(\mathbb{Q}_{p})$ in $\mathrm{Aut}(T)$.
By [14, Example 6.33] and [14, Corollary 6.28], $(G,T)$ is not (P)-closed (that is, following [14], $G$ does not coincide with its ($P_{1}$)-closure). However, it is well known that the relevant $\mathrm{PSL}_{2}(\mathbb{Q}_{p})$-action on $T$ is locally $\infty$-transitive (cf. [27, pp. 91 and 95]). In particular, $(G,T)$ is locally $\infty$-transitive.
Example 3.14 motivates the following question.
Question 3.15.
Is there a weakly locally $\infty$-transitive group action on a tree $(G,T)$ as in Convention 3.0.1 such that $G\leq\mathrm{Aut}(T)$ is closed, but $(G,T)$ is neither (P)-closed nor arc-transitive?